# Piotr Berman Meiram Murzabulatov

### v1

@@ 1.1 Our Results
As our starting point, we use an idea from [8] that gave an algorithm for approximating the (relative) distance to the nearest connected image with additive error ${\epsilon}$ with query $O(\frac{1}{{\epsilon}^{4}})$ and running time $\exp\left(O\left(\frac{1}{{\epsilon}}\right)\right)$. They observed that one can modify an image in a small number of pixels by drawing a grid on the image (as shown in Figures 1.2 and 1.2). In the resulting image, the distance to connectedness is determined by the properties of individual squares into which the grid lines partition the image.
@@
Definition 1.1 (Border connectedness).
A (sub)image $s$ is border-connected if for every black pixel $(i,j)$ of $s$, the image graph $G_{s}$ contains a path from $(i,j)$ to a pixel on the border of $s$. The property border connectedness, denoted ${\mathcal{C}}^{\prime}$, is the set of all border-connected images.
Our nonadaptive algorithm reads all pixels in each sampled square. Our adaptive algorithm further partitions each square into diamonds, as shown in Figure 3.2. It queries all pixels on the diamond lattice and then adaptively tries to catch a witness (that is, a small connected component) in one of the diamonds, using the lattice structure. (We could have partitioned into squares again, but partitioning into diamonds makes the proof cleaner and saves a constant factor in the analysis.)
@@
We also prove the first nontrivial lower bound for testing connectedness of images. Note that all nontrivial properties have query complexity $\Omega(1/{\epsilon})$, even for adaptive testers. In particular, for connectedness, this is the number of queries needed to distinguish between the white image and an image, where we color black a random subset of size ${\epsilon}n^{2}$ of the set of pixels with both coordinates divisible by 3. By standard arguments, it implies a lower bound of $\Omega(1/{\epsilon})$. Some properties of images, such is being a halfplane, can be tested nonadaptively (with 1-sided error) with $O(1/{\epsilon})$ queries [6]. We show that it is impossible for connectedness.
@@
Every nonadaptive (1-sided error) ${\epsilon}$-tester for connectedness of images must query $\Omega(\frac{1}{{\epsilon}}\log\frac{1}{{\epsilon}})$ pixels (for some family of images).
Every 1-sided error tester must catch a witness of disconnectedness in order to reject. This witness could include a connected component completely surrounded by white pixels. The difficulty for proving hardness is that, unlike in the case of finding a witness for disconnectedness of graphs, the algorithm does not have to read the whole connected component. Instead, it is sufficient to find a closed white loop with a black pixel inside it (and another black pixel outside it). As we discussed, it is sufficient for an algorithm to look for witnesses inside relatively small squares (specifically, squares with side length $O(1/{\epsilon})$), since adding a grid around such squares, as shown in Figure 1.2, would change $O({\epsilon}n^{2})$ pixels. But no matter how you had a witness inside such a square, it can be easily captured with $O(1/{\epsilon})$ queries if the border of the square is white.
@@ 1.2 Other Related Work
In addition to [12], connectedness testing and approximating the number of connected components in graphs in sublinear time was explored in [10, 5, 4]. Other property testing tasks studied in the pixel model of images, the model considered in this paper, include testing whether an image is a half-plane [19], convexity [19, 7, 6], and image partitioning properties [14]. Early implementations and applications to vision were provided in [14, 15, 16, 18]. Finally, general classes of matrix properties were investigated, including matrix-poset properties [11], earthmover resilient properties [2], hereditary properties [1], and classes of matrices that are free of specified patterns [3].
@@ 2 Definitions and Notation
We use $[0..{n})$ to denote the set of integers $\{0,1,\ldots,n-1\}$ and $[n]$ to denote $\{1,2,\ldots,n\}$. By $\log$ we mean the logarithm base 2.
@@
A property ${\mathcal{P}}$ is a set of images. The absolute distance from an image $M$ to a property ${\mathcal{P}}$, denoted $Dist(M,{\mathcal{P}}),$ is the smallest number of pixels in $M$ that need to be modified to get an image in ${\mathcal{P}}.$ The (relative) distance between an $n\times n$ image $M$ and a property ${\mathcal{P}}$ is $dist(M,{\mathcal{P}})=Dist(M,{\mathcal{P}})/n^{2}.$ We say that $M$ is ${\epsilon}$-far from ${\mathcal{P}}$ if $dist(M,{\mathcal{P}})\geq{\epsilon}$; otherwise, $M$ is ${\epsilon}$-close to ${\mathcal{P}}.$
@@ 3 Adaptive and Nonadaptive Property Testers for Connectedness
In this section, we present our testers for connectedness, proving Theorem 1.1. Both testers use the same top-level procedure, described in Algorithm 1. First, it samples random pixels to ensure that a black pixel is found. It will be used later to certify non-connectedness by producing a black pixel and an isolated black component. Then Algorithm 1 considers a logarithmic number of partitions of the image into subimages of the same size. For each partition, it samples a carefully selected number of these subimages and tests them for border connectedness (see Definition  1.1). This is where the two algorithms diverge. The nonadaptive algorithm tests for border connectedness using the subroutine Exhaustive-Square-Tester which queries all pixels in the sampled square and determines exactly if the square is border connected. The adaptive algorithm uses subroutine Diagonal-Square-Tester (Algorithm 2). If the top-level procedure finds a subimage that violates border connectedness and a black pixel outside that subimage, it rejects; otherwise, it accepts.
@@
For $i\in[0..{\log\frac{1}{{\epsilon}}})$, let $k_{i}=\frac{4}{{\epsilon}}\cdot 2^{-i}-1$. Call $(x,y)$ a grid pixel of level $i$ if $(k_{i}+1)|x$ or $(k_{i}+1)|y$.
For all coordinates $u,v$, which are divisible by $k_{i}+1$,
the $k_{i}\times k_{i}$ subimage that consists of pixels $[k_{i}]^{2}+(u,v)$ is called a square of level $i$. The set of all squares of level $i$ is denoted ${\cal{S}}_{i}$. Boundary pixels of a square of level $i$ are the pixels of the square which are adjacent to the grid pixels of level $i$. A square of any level that violates property ${\mathcal{C}}^{\prime}$ (see Definition 1.1) is called a witness.
@@
21 Query $\frac{4}{{\epsilon}}$ pixels uniformly at random with replacement. for $i=0$ to $\log\frac{1}{{\epsilon}}-1$ do
@@ 3 Repeat $2^{i+1}$ times:
    Run the border-connectedness subroutine with inputs $i,u,v$: if the tester is nonadaptive, use Exhaustive-Square-Tester; otherwise, use Diagonal-Square-Tester (Algorithm 2).
@@ 3.1 Effective Local Cost and the Structural Lemma
Claim 3.2 (Distance to Border Connectedness).
@@
If $s$ contains at most $\frac{k^{2}}{4}$ black pixels, we can make all of them white, i.e., modify at most $\frac{k^{2}}{4}$ pixels and obtain an image that satisfies ${\mathcal{C}}^{\prime}$. Now consider an image $s$ with more than $\frac{k^{2}}{4}$ black pixels, i.e., with less than $\frac{3k^{2}}{4}$ white pixels. Partition all pixels of $s$ into $3$ groups such that group $i\in\{0,1,2\}$ contains all pixels $(x,y)$, where $y\equiv i\pmod{3}$. Making all pixels of one group black produces an image that satisfies ${\mathcal{C}}^{\prime}$. By averaging, at least one group has less than $\frac{k^{2}}{4}$ white pixels. Making all these white pixels black results in an image that satisfies ${\mathcal{C}}^{\prime}$. This completes the proof. ∎
@@
Fix a level $i$ and a square $s\in\mathcal{S}_{i}$. Let $desc(s,j)$ denote the set of all squares of level $j\geq i$ inside $s$. (In particular, $desc(s,i)$ contains only $s$.) We will prove by induction that for all integer $j\in[i,\log\frac{1}{{\epsilon}}-1),$
@@
$\sum\nolimits_{q\in desc(s,m+1)}lc(q)\leq\sum\nolimits_{q\in desc(s,m+1)}elc(q)+\sum\nolimits_{q\in desc(s,m+2)}lc(q)$
and
@@
where the final equality holds because in every square of level $i=\log\frac{1}{{\epsilon}}-1$, we have $k_{i}=7$, and consequently, by 3.2, the local cost is at most $\frac{7^{2}}{4}<2\cdot 7$, i.e., it is equal to the effective local cost of that square. Summing up (2) for all squares in $\mathcal{S}_{0}$, we get
@@
where the last equality is obtained by switching the order of summations and rearranging the second summation in terms of levels. ∎
@@ 3.2 Testing Border-Connectedness
In this section, we state and analyze our adaptive border connectedness subroutine after defining the concepts used in it. We state the guarantees of both border connectedness subroutines in Lemma 3.4.
For the adaptive subroutine, we partition the square into diamonds and fences surrounding them, as described in Definition 3.3. The subroutine queries all pixels on the fences and categorizes diamonds into those whose black pixels are potentially connected to the border (set $B$ in Algorithm 2) and those whose black pixels are definitely not (set $A$ in Algorithm 2). Then it tries to find a black pixel in a diamond from set $A$ or (using BFS) an isolated black component in diamonds from set $B$. Observe that either of the two provides evidence that the square is not border-connected.
Figure 3.2: An example of execution of Algorithm 2. Black lines represent lattice pixels. The blue diamonds are included in $B$ because they are adjacent to the border of the square. The purple dots represent black pixels in the lattice. The green diamonds are added to $B$ during the BFS because their fences contain black pixels. The white diamonds remain in $A$.
@@
Note that lattice pixels are not part of any diamond. Moreover, some diamonds are partial (those that have pixels from the border of the square.)
input : parameters $i$, $u$ and $v$; access to an $n\times n$ matrix $M$.
4321 Let $s$ be a square of level $i$ that consists of pixels $[k_{i}]^{2}+(u,v)$ and $m_{i}$ be the largest odd integer less than or equal to $\lceil\sqrt{k_{i}/\log k_{i}}\rceil$. Query all the diagonal lattice pixels of $s$ (see Definition 3.3). Initialize $B$ to be the set of all diamonds of $s$ that contain a border pixel of $s$. Initialize $A$ to be the set of the remaining diamonds of $s.$ While $\exists d_{1}\in B$ and $\exists d_{2}\in A$ such that $d_{1}$ and $d_{2}$ have a black pixel in the common portion of their fences, move $d_{2}$ from $A$ to $B$. Repeat $\frac{1}{2}k_{i}m_{i}$ times:
@@
  • (b)
    If $p$ is a black pixel from a diamond in $B$, pick a natural number $x\in[k_{i}^{2}]$ from the distribution with the probability mass function $f(j)=\frac{1}{j(j+1)}$ for all $j\in[k^{2}_{i}-1]$ and $f(j)=\frac{1}{j}$ for $j=k^{2}_{i}$. (Observe that $\Pr(x\geq j)=\frac{1}{j}$ for all $j\in[k_{i}^{2}]$.)
  • 5 Starting from $p$, perform a BFS of its connected component, halting if $x+1$ black pixels in $p$’s component are discovered or a black pixel on the border of square $s$ is reached. If the BFS halts after discovering at most $x$ black pixels, none of which are on the border of $s,$ reject. Accept.
@@
    Now we prove the statement for Diagonal-Square-Tester. Let $A$ and $B$ be defined as in Algorithm 2 after Step 3. Let $a$ be the number of black pixels in all the diamonds of the set $A$ and on the fences of those diamonds. Let $b$ be the number of connected components of the image graph that are formed by black pixels in all the diamonds in the set $B$ and in their fences and that contain no pixels on the border of $s$. Next, we prove that
    $bm_{i}+a\geq lc(s)\geq elc(s).$ (3)
    The second inequality in (3) holds by Definition 3.2. To prove the first, we will show how to obtain a square in ${\mathcal{C}}^{\prime}$ by modifying at most $bm_{i}+a$ pixels in $s$.
    We claim that we can connect all $b$ connected components to each other and to the border of the square by modifying at most $bm_{i}$ pixels. To see this, notice that any two black pixels in the same diamond can be connected to each other by changing less than $m_{i}$ pixels (by taking any Manhattan-distance shortest path between the two pixels and making it all black). To prove the claim, we can connect the $b$ connected components to the border of the square in the order their diamonds were added to $B$ by the algorithm. The initial diamonds placed in $B$ have at least one pixel on the border of the square, so their connected components can be directly connected to the border (using at most $m_{i}$ pixels per connected component). Now assume that we already connected to the border all components that have pixels in the diamonds added to $B$ so far. When the algorithm moves some diamond $d_{2}$ from $A$ to $B$, it is done because there is already a diamond $d_{1}$ in $B$ such that $d_{1}$ and $d_{2}$ have a black pixel $\beta$ in the common portion of their fences. Then $\beta$ must be already connected to the border. If there are any connected components in $d_{2}$ that are not connected to the border yet, we can fix that by connecting them to $\beta$ (using at most $m_{i}$ pixels per connected component). We proceed like this until all $b$ connected components of the diamonds that were added to $B$ by the algorithm are connected to the border of the square. At this point, we have changed at most $bm_{i}$ pixels.
    Recall that we have $a$ black pixels in all the diamonds of the set $A$ and on the fences of those diamonds. We make all of them white. After these at most $bm_{i}+a$ modifications to $s$, we obtain a square in ${\mathcal{C}}^{\prime}$. Thus, (3) holds.
    Observe that for $x\in[0,1]$,
@@
    Consider one iteration of Step 2 of Algorithm 2. Let $E_{1}$ be the event that $p$ is one of the $a$ black pixels in the diamonds of the set $A$. Let $CC$ denote the set of the $b$ connected components from the diamonds of the set $B$. Let $E_{2}$ be the event that $p$ is in one of the connected components in $CC$ and that its connected component is completely discovered by the BFS in Step 2b. Then the probability that one iteration of Step 2 of Algorithm 2 rejects square $s$ is equal to $\Pr[E_{1}]+\Pr[E_{2}]$. Since $p$ is chosen uniformly from a square of size $k_{i}\times k_{i}$,
    $\Pr[E_{1}]=\frac{a}{k^{2}_{i}}.$
@@
    $\Pr[E_{2}]=\sum_{C\in CC}\Pr_{p}[p\in C]\cdot\Pr_{x}[x\geq|C|]=\sum_{C\in CC}\frac{|C|}{k^{2}_{i}}\cdot\frac{1}{|C|}=\sum_{C\in CC}\frac{1}{k^{2}_{i}}=\frac{b}{k^{2}_{i}}.$
    Then, by (3) and (4) and since $\frac{elc(s)}{2k_{i}}\leq 1$ by Definition 3.2, the probability that all $\frac{1}{2}k_{i}m_{i}$ iterations of Step 2 of Algorithm 2 proceed without rejecting is
    $\left(1-\frac{a+b}{k^{2}_{i}}\right)^{\frac{1}{2}k_{i}m_{i}}\leq e^{-\frac{a+b}{2k_{i}}\cdot m_{i}}\leq e^{-\frac{a+bm_{i}}{2k_{i}}}\leq e^{{-\frac{elc(s)}{2k_{i}}}}\leq 1-\frac{elc(s)}{2k_{i}}(1-e^{-1}).$
@@
    Let $E$ be the event that,
    in this iteration, the subroutine in Step 1b rejects. Since events $E_{s}$ are disjoint and $E=\bigcup_{s\in\mathcal{S}_{i}}E_{s},$

### v2

@@ 1.1 Our Results
As our starting point, we use an idea from [7]. That work gave an algorithm for approximating the (relative) distance to the nearest connected image within additive error ${\epsilon}$ using $O(\frac{1}{{\epsilon}^{4}})$ queries in $\exp\left(O\left(\frac{1}{{\epsilon}}\right)\right)$ time. The authors observed that one can modify an image in a small number of pixels by drawing a grid on the image (as shown in Figures 1.2 and 1.2). In the resulting image, the distance to connectedness is determined by the properties of individual squares into which the grid lines partition the image.
@@
Definition 1.1 (Border-connectedness).
A (sub)image $s$ is border-connected if for every black pixel $(i,j)$ of $s$, the image graph $G_{s}$ contains a path from $(i,j)$ to a pixel on the border of $s$. The property border-connectedness, denoted ${\mathcal{C}}^{\prime}$, is the set of all border-connected images.
Our nonadaptive algorithm reads all pixels in each sampled square. Our adaptive algorithm further partitions each square into diamonds, as shown in Figure 3.2. (We could have partitioned into squares again, but partitioning into diamonds makes the proof cleaner and saves a constant factor in the analysis.) The algorithm queries all pixels on the diamond lattice and then adaptively tries to catch a witness (that is, a small connected component) in one of the diamonds, using the lattice structure. It partitions the diamonds into those that can be easily connected to the border and those that cannot. It tries to catch either a black pixel in the diamond in the latter group or, alternatively, a small connected component in the diamond in the former group by running a breadth-first search (BFS) for a random number of steps, drawn from a carefully selected distribution.
Even though we use the notion of border-connectedness from [7], our algorithms and analysis are different. The authors of [7] approximate the distance to connectectedness by partitioning the image into squares with the side length $\Theta(1/{\epsilon})$, sampling such a square, and then returning the exact distance to border-connectedness of the sampled square. (Computing this distance exactly, even in exponential time, is highly nontrivial.) In contrast, we use multiple partitions of the image. Moreover, our adaptive subroutine queries a small fraction of the pixels in the sampled square to certify that the square is not border-connected. The notions we use in the analysis of our algorithms and the accompanying correction procedures are also new.
@@
We also prove the first nontrivial lower bound for testing connectedness of images. Note that all nontrivial properties have query complexity $\Omega(1/{\epsilon})$, even for adaptive testers. In particular, for connectedness, this is the number of queries needed to distinguish between the white image and an image, where we color black a random subset of size ${\epsilon}n^{2}$ of the set of pixels with both coordinates divisible by 3. By standard arguments, it implies a lower bound of $\Omega(1/{\epsilon})$. Some properties of images, such as being a halfplane, can be tested nonadaptively (with 1-sided error) with $O(1/{\epsilon})$ queries [5]. We show that it is impossible for connectedness.
@@
Every nonadaptive, 1-sided error ${\epsilon}$-tester for connectedness of images must query $\Omega(\frac{1}{{\epsilon}}\log\frac{1}{{\epsilon}})$ pixels (for some family of images).
Every 1-sided error tester must catch a witness of disconnectedness in order to reject. This witness could include a connected component completely surrounded by white pixels. The difficulty for proving hardness is that, unlike in the case of finding a witness for disconnectedness of graphs, the algorithm does not have to read the entire connected component. Instead, it is sufficient to find a closed white loop with a black pixel inside it (and another black pixel outside it). As we discussed, it is enough for an algorithm to look for witnesses inside relatively small squares (specifically, squares with side length $O(1/{\epsilon})$), since adding a grid around such squares, as shown in Figure 1.2, would change $O({\epsilon}n^{2})$ pixels. But no matter how a witness inside such a square looks like, it can be easily captured with $O(1/{\epsilon})$ queries if the border of the square is white.
@@ 1.2 Other Related Work
In addition to [13], connectedness testing and approximating the number of connected components in graphs in sublinear time was explored in [10, 9, 3]. Other property testing tasks studied in the pixel model of images, the model considered in this paper, include testing whether an image is a half-plane [20, 21, 18], convexity [20, 21, 18, 6, 5], and image partitioning properties [14]. Early implementations and applications to vision were provided in [14, 15, 16, 19]. Finally, general classes of matrix properties were investigated, including matrix-poset properties [11], earthmover resilient properties [2], hereditary properties [1], and classes of matrices that are free of specified patterns [4].
@@ 2 Definitions and Notation
We use $[0..{n})$ to denote the set of integers $\{0,1,\ldots,n-1\}$ and $[n]$ to denote $\{1,2,\ldots,n\}$. By $\log$ we mean the logarithm base 2. For a set $S\subset[0..{n})^{2}$ and $(i,j)\in[0..{n})^{2}$, we define $S+(i,j)=\{(x+i,y+j):\leavevmode\nobreak\ (x,y)\in S\}$.
@@
A property ${\mathcal{P}}$ is a set of images. We say that an image $M$ satisfies ${\mathcal{P}}$ if $M\in{\mathcal{P}}$; otherwise, $M$ violates ${\mathcal{P}}.$ The absolute distance from an image $M$ to a property ${\mathcal{P}}$, denoted $Dist(M,{\mathcal{P}}),$ is the smallest number of pixels in $M$ that need to be modified to get an image in ${\mathcal{P}}.$ The (relative) distance between an $n\times n$ image $M$ and a property ${\mathcal{P}}$ is $dist(M,{\mathcal{P}})=Dist(M,{\mathcal{P}})/n^{2}.$ We say that $M$ is ${\epsilon}$-far from ${\mathcal{P}}$ if $dist(M,{\mathcal{P}})\geq{\epsilon}$; otherwise, $M$ is ${\epsilon}$-close to ${\mathcal{P}}.$
@@ 3 Adaptive and Nonadaptive Property Testers for Connectedness
In this section, we present our testers for connectedness, proving Theorem 1.1. Both testers use the same top-level procedure, described in Algorithm 1. First, it samples random pixels to ensure that a black pixel is found. The black pixel will be used later together with an isolated black component to certify non-connectedness. Then Algorithm 1 considers a logarithmic number of partitions of the image into subimages of the same size. For each partition, it samples a carefully selected number of these subimages and tests them for border-connectedness (see Definition  1.1). This is where the two algorithms diverge. The nonadaptive algorithm tests for border-connectedness using the subroutine Exhaustive-Square-Tester which queries all pixels in the sampled square and determines exactly if the square is border-connected. This can be done by a BFS on the image graph for the sampled square that checks if all the black pixels can be reached from the border. The adaptive algorithm uses subroutine Diagonal-Square-Tester (Algorithm 2). If the top-level procedure finds a subimage that violates border-connectedness and a black pixel outside that subimage, it rejects; otherwise, it accepts.
@@
For $i\in[0..{\log\frac{1}{{\epsilon}}})$, let $k_{i}=\frac{4}{{\epsilon}}\cdot 2^{-i}-1$. Call $(x,y)$ a grid pixel of level $i$ if $(k_{i}+1)|x$ or $(k_{i}+1)|y$. For all coordinates $u,v$, which are divisible by $k_{i}+1$, the $k_{i}\times k_{i}$ subimage that consists of pixels $[k_{i}]^{2}+(u,v)$ is called a square of level $i$. The set of all squares of level $i$ is denoted ${\cal{S}}_{i}$. Boundary pixels of a square of level $i$ are the pixels of the square which are adjacent to the grid pixels of level $i$. A square of any level that violates property ${\mathcal{C}}^{\prime}$ (see Definition 1.1) is called a witness.
@@
1 Query $\frac{8}{{\epsilon}}$ pixels uniformly at random with replacement.
2 for $i=0$ to $\log\frac{1}{{\epsilon}}-1$ do
@@ 3 Repeat $2^{i+1}$ times:
    Run the border-connectedness subroutine on the square $s$: if the tester is nonadaptive, use Exhaustive-Square-Tester (see the beginning of Section 3); otherwise, use Diagonal-Square-Tester (Algorithm 2) with inputs $i,u,v$.
@@ 3.1 Effective Local Cost and the Structural Lemma
Claim 3.2 (Distance to border-connectedness).
@@
If $s$ contains at most $\frac{k^{2}}{4}$ black pixels, we can make all of them white, i.e., modify at most $\frac{k^{2}}{4}$ pixels and obtain an image that satisfies ${\mathcal{C}}^{\prime}$. Now consider an image $s$ with more than $\frac{k^{2}}{4}$ black pixels, i.e., with less than $\frac{3k^{2}}{4}$ white pixels. Partition all pixels of $s$ into $3$ groups such that group $i\in\{0,1,2\}$ contains all pixels $(x,y)$, where $y\equiv i\pmod{3}$. Making all pixels of one group black produces an image where every third row is completely black. Note that completely black rows are also connected to the border. Every black pixel in the recolored image that is not in such a row is either directly above or directly below a black row and, consequently, also connected to the border. Thus, the new image satisfies ${\mathcal{C}}^{\prime}$. By averaging, at least one group has less than $\frac{k^{2}}{4}$ white pixels. Making all these white pixels black results in an image that satisfies ${\mathcal{C}}^{\prime}$, proving the claim. ∎
@@
Fix a level $i$ and a square $s\in\mathcal{S}_{i}$. For a level $j\geq i$, let $desc(s,j)$ denote the set of all squares of level $j$ inside $s$. (In particular, $desc(s,i)$ contains only $s$.) We will prove by induction that for every integer $j\in[i,\log\frac{1}{{\epsilon}}-1),$
@@
$\sum\nolimits_{q\in desc(s,m+1)}lc(q)\leq\sum\nolimits_{q\in desc(s,m+1)}elc(q)+\sum\nolimits_{q\in desc(s,m+2)}lc(q).$
By the induction hypothesis,
@@
where the final equality holds because in each square of level $i=\log\frac{1}{{\epsilon}}-1$, we have $k_{i}=\frac{4}{{\epsilon}}\cdot 2^{-i}-1=7$, and consequently, by 3.2, the local cost is at most $\frac{7^{2}}{4}<2\cdot 7$, i.e., it is equal to the effective local cost of that square. Summing up (2) for all squares in $\mathcal{S}_{0}$, we get
@@
where the last equality is obtained by switching the order of summations and rearranging the second summation in terms of the levels. ∎
@@ 3.2 Testing Border-Connectedness
In this section, we state and analyze our adaptive border-connectedness subroutine after defining the concepts used in it. We state the guarantees of both border-connectedness subroutines in Lemma 3.4.
For the adaptive subroutine, we partition the square into diamonds and fences surrounding them, as described in Definition 3.3. The subroutine queries all pixels on the fences and categorizes diamonds into those whose black pixels are potentially connected to the border (set $B$ in Algorithm 2) and those whose black pixels are definitely not (set $A$ in Algorithm 2). Then it tries to find a black pixel in a diamond from set $A$ or (using BFS) an isolated black component in diamonds from set $B$. Observe that either of the two provides evidence that the square is not border-connected. In order to ensure that Algorithm 2 does not get stuck investigating a large connected component, BFS is stopped after a random number of steps, $x$. This idea was used for estimating the number of connected components in a graph by Berenbrink, Krayenhoff, and Mallmann-Trenn [3]. However, they select $x$ from a different distribution: specifically, with $\Pr[x\geq j]=1/j^{2}$, whereas we set $\Pr[x\geq j]=1/j$.
Figure 3.2: An example of execution of Algorithm 2. Black lines represent lattice pixels. The blue diamonds are included in $B$ because they contain pixels from the border of the square. The purple dots represent black pixels in the lattice. The green diamonds are added to $B$ during the BFS because their fences contain black pixels. The white diamonds remain in $A$.
@@
Observe that lattice pixels are not part of any diamond. Moreover, some diamonds (those that have pixels from the border of the square) are partial.
input : level $i$, coordinates $u,v\in[n]$; access to an $n\times n$ matrix $M$.
Let $s$ be a square of level $i$ that consists of pixels $[k_{i}]^{2}+(u,v)$ and $m_{i}$ be the largest odd integer less than or equal to $\lceil\sqrt{k_{i}/\log k_{i}}\rceil$.
1 Query all the diagonal lattice pixels of $s$ (see Definition 3.3).
2 Initialize $B$ to be the set of all diamonds of $s$ that contain a border pixel of $s$. Initialize $A$ to be the set of the remaining diamonds of $s.$
3 While $\exists d_{1}\in B$ and $\exists d_{2}\in A$ such that $d_{1}$ and $d_{2}$ have a black pixel in the common portion of their fences, move $d_{2}$ from $A$ to $B$.
4 Repeat $\frac{1}{2}k_{i}m_{i}$ times:
@@
  3. (c)
    If $p$ is a black pixel from a diamond in $B$, pick a natural number $x\in[k_{i}^{2}]$ from the distribution with the probability mass function $f(j)=\frac{1}{j(j+1)}$ for all $j\in[k^{2}_{i}-1]$ and $f(j)=\frac{1}{j}$ for $j=k^{2}_{i}$. (Observe that $\Pr[x\geq j]=\frac{1}{j}$ for all $j\in[k_{i}^{2}]$.) Starting from $p$, perform a BFS of its connected component, halting if $x+1$ black pixels in $p$’s component are discovered or a black pixel on the border of square $s$ is reached. If the BFS halts after discovering at most $x$ black pixels, none of which are on the border of $s,$ reject.
5 Accept.
@@
Note that Algorithm 2 could reject right after Step 2 if any diamond in $A$ has a black pixel on its fence. However, we skipped this operation to keep the algorithm concise, since it is not needed to provide the guarantees of Theorem 1.1.
The success probability of Algorithm 2 is analyzed in Lemma 3.4. The intuition behind the analysis is that every square $s$ that has a high local cost (i.e., is far from border-connectedness) has either many black pixels in the diamonds of $A$ or many black connected-components in the diamonds of $B$. This is because we can “fix” $s$ by making all black pixels in $A$ white and by connecting all black connected components of $B$ to the border at the cost of recoloring at most $m_{i}$ white pixels per connected component.
@@
Now we prove the statement for Diagonal-Square-Tester. Let $A$ and $B$ be defined as in Algorithm 2 after Step 3. Let $\mathbf{a}$ be the number of black pixels in all the diamonds of the set $A$ and on the fences of these diamonds. Consider the connected components of the image graph that are formed by black pixels in all the diamonds in the set $B$ and in their fences. Let $\cal{K}$ be the set of all such connected components that contain no pixels on the border of $s$, and define $\mathbf{b}=|\cal{K}|.$
Next, we prove that
$\mathbf{b}m_{i}+\mathbf{a}\geq lc(s)\geq elc(s).$ (3)
The second inequality in (3) holds by Definition 3.2. To prove the first, we will show how to obtain a square in ${\mathcal{C}}^{\prime}$ by modifying at most $\mathbf{b}m_{i}+\mathbf{a}$ pixels in $s$.
We claim that we can connect all $\mathbf{b}$ connected components to each other and to the border of the square by modifying at most $\mathbf{b}m_{i}$ pixels. To see this, notice that any two black pixels in the same diamond can be connected to each other by changing less than $m_{i}$ pixels (by taking any Manhattan-distance shortest path between the two pixels and making it all black). To prove the claim, we can connect the $\mathbf{b}$ connected components to the border of the square in the order their diamonds were added to $B$ by the algorithm. The initial diamonds placed in $B$ have at least one pixel on the border of the square, so their connected components can be directly connected to the border (using at most $m_{i}$ pixels per connected component). Now assume that we already connected to the border all components that have pixels in the diamonds added to $B$ so far. When the algorithm moves some diamond $d_{2}$ from $A$ to $B$, it is done because there is already a diamond $d_{1}$ in $B$ such that $d_{1}$ and $d_{2}$ have a black pixel $\beta$ in the common portion of their fences. Then $\beta$ must be already connected to the border. If there are any connected components in $d_{2}$ that are not connected to the border yet, we can fix that by connecting them to $\beta$ (using at most $m_{i}$ pixels per connected component). We proceed like this until all $\mathbf{b}$ connected components of the diamonds that were added to $B$ by the algorithm are connected to the border of the square. At this point, we have changed at most $\mathbf{b}m_{i}$ pixels.
Finally, we color white all the $\mathbf{a}$ black pixels in the diamonds of the set $A$ and on their fences. Altogether, after these at most $\mathbf{b}m_{i}+\mathbf{a}$ modifications to $s$, we obtain a square in ${\mathcal{C}}^{\prime}$. Thus, (3) holds.
Next, we use (3) to analyze the success probability of Algorithm 2. Observe that for $x\in[0,1]$,
@@
Consider one iteration of Step 2 of Algorithm 2. Recall that $p$ is the pixel selected in that step. Let $E_{1}$ be the event that $p$ is one of the $\mathbf{a}$ black pixels in the diamonds of the set $A$. Let $E_{2}$ be the event that $p$ is in one of the connected components in $\cal{K}$ and that its connected component is completely discovered by the BFS in Step 2c. Then the probability that one iteration of Step 2 of Algorithm 2 rejects square $s$ is equal to $\Pr[E_{1}]+\Pr[E_{2}]$. Since $p$ is chosen uniformly from a square of size $k_{i}\times k_{i}$,
$\Pr[E_{1}]=\frac{\mathbf{a}}{k^{2}_{i}}.$
@@
$\Pr[E_{2}]=\sum_{C\in\cal{K}}\Pr_{p}[p\in C]\cdot\Pr_{x}\big{[}x\geq|C|\big{]}=\sum_{C\in\cal{K}}\frac{|C|}{k^{2}_{i}}\cdot\frac{1}{|C|}=\sum_{C\in\cal{K}}\frac{1}{k^{2}_{i}}=\frac{\mathbf{b}}{k^{2}_{i}}.$
Then, by (3) and (4) and since $\frac{elc(s)}{2k_{i}}\leq 1$ by Definition 3.2, the probability that all $\frac{1}{2}k_{i}m_{i}$ iterations of Step 2 of Algorithm 2 complete without rejecting is
$\left(1-\frac{\mathbf{a}+\mathbf{b}}{k^{2}_{i}}\right)^{\frac{1}{2}k_{i}m_{i}}\leq e^{-\frac{\mathbf{a}+\mathbf{b}}{2k_{i}}\cdot m_{i}}\leq e^{-\frac{\mathbf{a}+\mathbf{b}m_{i}}{2k_{i}}}\leq e^{{-\frac{elc(s)}{2k_{i}}}}\leq 1-\frac{elc(s)}{2k_{i}}(1-e^{-1}).$
@@
Let $E$ be the event that, in this iteration, the subroutine in Step 1b rejects. Since events $E_{s}$ are disjoint and $E=\bigcup_{s\in\mathcal{S}_{i}}E_{s},$